OpenStudy (anonymous):

limits question! \[\lim_{t \rightarrow 0} \left[ 1\div t \sqrt{1+t} -1\div t \right]\]

hero (hero):

rationalize

OpenStudy (anonymous):

find a common denominator and things should cancel

OpenStudy (anonymous):

thats two separate fractions right?

hero (hero):

yup

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

find a common denominator the usual way and it should work out

OpenStudy (anonymous):

please show me steps.

hero (hero):

I'm sleepy. I'm going to bed. QED will show you hopefully

OpenStudy (anonymous):

i will

OpenStudy (anonymous):

ok, gnight. QED, thank you. :)

OpenStudy (anonymous):

last one tho bcuz this is annoying to type out

OpenStudy (anonymous):

ok, thank you so much for all the help tho.

OpenStudy (anonymous):

you truly are a lifesaver.

OpenStudy (anonymous):

\[=(1-\sqrt{1+t})/t \sqrt{1+t} after combining terms\]

OpenStudy (anonymous):

then you rationalize \[((1-\sqrt{1+t})/t \sqrt{1+t}) * (1+\sqrt{1+t})/(1+\sqrt{1+t}))\]

OpenStudy (anonymous):

then u have only t left on top, which cancels the t on the bottom (DO NO MULTIPLY BOTTOM OUT)

OpenStudy (anonymous):

then just plug in 0 and youre good should get.... 1/2

hero (hero):

frac{a}{b}

OpenStudy (anonymous):

thanks hero

OpenStudy (anonymous):

sorry, -1/2

hero (hero):

It's a little tricky posting those with square roots, but you'll figure it out

OpenStudy (anonymous):

ok, thank you both

OpenStudy (anonymous):

np

hero (hero):

I didn't even do anything

OpenStudy (anonymous):

well you tried to help. so ty, that counts. :)